What volume (in mL) of 3.75×10^−2 M HNO3 is needed to titrate 45.0 mL of 1.15×10^−2 M LiOH to the equivalence point?
Answer #1

Balanced chemical equation is:
LiOH + HNO3 ---> LiNO3 + H2O


Here:
M(LiOH)=0.0115 M
M(HNO3)=0.0375 M
V(LiOH)=45.0 mL

According to balanced reaction:
1*number of mol of LiOH =1*number of mol of HNO3
1*M(LiOH)*V(LiOH) =1*M(HNO3)*V(HNO3)
1*0.0115 M *45.0 mL = 1*0.0375M *V(HNO3)
V(HNO3) = 13.8 mL

Answer: 13.8 mL